PURE ROTATIONAL SPECTRUM AND STRUCTURE OF ZIRCONIUM DIOXIDE

\underline{DALE J. BRUGH} and RICHARD D. SUENRAM, National Institute of Standards and Technology, Optical Technology Division, Gaithersburg, MD 20899-8441; WALTER J. STEVENS, National Institute of Standards and Technology, Physical and Chemical Properties Division, Gaithersburg, MD 20899-8380.

The pure rotational spectrum of the asymmetric top ZrO_{2} has been collected using a Fourier transform microwave spectrometer that employed a laser ablation molecular beam source. Four rotational transitions for each of five Zr^{16}O_{2} isotopomers have been recorded. The rotational constants of the ^{90}Zr^{16}O_{2} isotope were determined to be A = 19881.352 ± 0.068 MHz, B = 7693.895 ± 0.021 MHz, and C = 5533.111 ± 0.036 MHz. The r_{0} structure was determined to possess a Zr-O bond length of 1.7710 ± 0.0007 Å, and an O-Zr-O bond angle of 108.11 ± 0.08^{∘}. The electric dipole moment has been measured for the ^{90}Zr^{16}O_{2} isotope and found to be μ_{b} = 7.80 ± 0.02 Debye. The nuclear quadrupole hyperfine structure for the ^{91}Zr^{16}O_{2} isotope has also been recorded and analyzed, yielding χ_{aa} = 115.94 ± 0.16 MHz, χ_{bb} = -37.55 ± 0.33 MHz, and χ_{cc} = -78.39 ± 0.16 MHz. High-level density functional theory calculations have yielded a structure that agrees well with the values determined experimentally.